### v1

@@ Abstract
Monitoring maternal and fetal health during pregnancy is crucial for preventing adverse outcomes. While tests such as ultrasound scans offer high accuracy, they can be costly and inconvenient. Telehealth and more accessible body shape information provide pregnant women with a convenient way to monitor their health. This study explores the potential of 3D body scan data, captured during the 18-24 gestational weeks, to predict adverse pregnancy outcomes and estimate clinical parameters. We developed a novel algorithm with two parallel streams which are used for extract body shape features: one for supervised learning to extract sequential abdominal circumference information, and another for unsupervised learning to extract global shape descriptors, alongside a branch for demographic data. Our results indicate that 3D body shape can assist in predicting preterm labor, gestational diabetes mellitus (GDM), gestational hypertension (GH), and in estimating fetal weight. Compared to other machine learning models, our algorithm achieved the best performance, with prediction accuracies exceeding 88% and fetal weight estimation accuracy of 76.74% within a 10% error margin, outperforming conventional anthropometric methods by 22.22%.
@@
Pregnancy outcomes, 3D body scan, Machine learning
@@ 1 Introduction
Although telehealth offers convenience, it has disadvantage of lacking in laboratory tests [10, 11]. To address this issue, it is crucial to explore biomarkers which can be easily obtained through widely-available devices for health monitoring. The 3D body shape is an ideal data modality that can be collected in non-invasive ways, especially in the field of obstetrics, where distinct body shape changes occur on pregnant women. Meanwhile, previous studies have shown a strong correlation between maternal anthropometric measurements and the health status of mothers and fetuses [12, 13, 14]. Ay et al. found that maternal body mass index (BMI) during pregnancy positively correlates with the fetal weight. Maternal height, pre-pregnancy BMI, and gestational weight gain are also found to correlate with potential risks of having a small or large for gestational age child [15]. Risk of preeclampsia and cesarean delivery have been associated with maternal weight, height and body circumferences [16, 17]. These promising results indicate the potential for using body shape features as additional resources in telehealth for pregnancy. Moreover, comparing to traditional anthropometric features, detailed 3D body scan data can provide more body shape information, such as the shape, volume and position of gravid uterus.
@@
Currently, 3D body scanning technology in the field of obstetrics and gynecology is still in its early stages. However, a growing number of research has demonstrated the efficacy of assessing health conditions by analyzing 3D body shapes. Machine learning algorithms, in particular, have shown high potential in extracting latent information from 3D body shapes. For example, Lu et al. proposed a method to predict body fat percentage from 3D body scans using machine learning algorithms [20]. Zheng et al. designed a new shape descriptor and a neural network to predict appendicular skeletal muscle mass [21]. Wang et al. designed a bi-channel network to detect hepatic steatosis using 3D body shape reconstructed from CT iso-surfaces [22]. These studies present the possibility of employing 3D body shape in the field of obstetrics.
In this study, we aim to apply 3D body shape analysis in maternal and fetal health status assessment. Our main contributions are summarized as follows:
@@
    We used 3D body scans and basic demographic information to predict the risk of preterm labor, GDM, GH, and the likelihood of delivery by cesarean section. We also used the same body shape features to estimate fetal weight and MVP. To the best of our knowledge, this is the first study that uses 3D body scans of pregnant women to assess maternal and fetal health status.
@@
    We conducted experiments comparing our algorithm using 3D body scans as inputs with the baseline method using support vector machine (SVM) and anthropomorphic measurements. Additionally, we also compared our algorithm with other well-performing machine learning algorithms using the same 3D body shape inputs. These experimental results not only indicate the feasibility of using 3D body scan to assess maternal and fetal health status, but also demonstrate the accuracy and efficiency of our new approach.
@@ 2.1 Association between anthropometric measurement and pregnancy health status
To simply predict the likelihood of adverse pregnancy outcomes, researchers explored the relationships between maternal anthropometric parameters and a range of adverse outcomes, and they found a strong correlation between them. Boucher et al. conducted a study which revealed associations between the probability of cesarean delivery and anthropometric measurements such as weight, BMI, waist circumference and skinfold thickness [17]. Sina et al. found that increasing BMI, weight, waist circumference, and waist-to-height ratio were associated with an increased risk of GDM, with BMI and waist-to-height ratio showing stronger correlations [23]. Ebrahimi-Mameghani et al. concluded that early pregnancy BMI and waist circumference were associated with the risk of GH, preeclampsia, and preterm labor [24].
@@
Anthropometric techniques have been widely used to capture a small part of body shape features for a long time, but these methods cannot fully represent the whole body shape [32, 33]. Moreover, although these methods do not require complex equipment, they still need to be carried out by professionals.
@@
Recently, some researchers investigated the applications of 3D body scanning in obstetrics. Glinkowski et al. used 3D surface topography method to reveal postural changes of pregnant women and studied the relationship between spinal curvatures and low back pain during pregnancy [34]. Dathan-Stumpf et al. captured anthropometric parameters from 3D scan models of pregnant women to predict successful vaginal breech deliveries. The results showed that the prediction accuracy of the method using 3D scanning measurements is at least as good as Magnetic Resonance Imaging (MRI) diagnosis [35]. These studies further indicated the usability of 3D body scans in obstetrics. However, they are still limited by the extraction of a limited number of anthropometric measurements from 3D models rather than learning information directly from the 3D models as a new imaging modality.
@@
To fully leverage information from the complicated representations of 3D models, which are often comprised of thousands or millions of polygons or 3D points [36], researchers are investigating more efficient body representations and methodologies and evaluating them on different medical applications. Xie et al. proposed to use frontal whole-body silhouettes for estimating body composition. Test results on estimating fat mass index and fat-free mass index supported the usability of this silhouette-based method [37]. Ng et al. proposed a method to estimate regional body composition by using regional circumferences, areas, and volumes obtained from 3D body models with a linear model [18]. Lu et al. used level circumferences to represent 3D body shapes. With the extracted level circumferences, they employed a Bayesian network to predict pixel-level body composition and body fat percentage for the 2D projection of 3D body [38]. Wang et al. constructed a body shape descriptor called Shape Map from body shape contour of CT scan slices and used it to predict whole-body fat percentage and visceral fat percentage [39]. Zheng et al. used abdominal level circumferences sampled from 3D body shapes to assess hepatic steatosis [40]. Roy et al. obtained the trajectory of the vertebral column from body shape and derived potential lateral deviations of the spine and rotation of the vertebrae. This method performed well in the estimations of the lateral deviation of the spine for mild and moderate scoliosis [41]. Su et al. explored the diagnosis of type II diabetes using features selected from 3D scan data through machine learning algorithms like backpropagation neural networks and decision trees [42]. Although these studies are not specifically conducted in obstetrics field, the various methods explored to represent 3D body shape and extract body shape features are worth learning.
@@ 3.1 Dataset
Our dataset comprises a total of 60 study participants, who were recruited at the George Washington University Medical Faculty Associates. The demographic characteristics of this study population is listed in Table 1. To mitigate the influence of confounding factors, we implemented the recruitment following these exclusion criteria: (1) being under the age of 18; (2) having carried multiple gestations; (3) having been diagnosed with an enlarged fibroid uterus; (4) having a BMI exceeding 60; (5) having any unstable medical or emotional condition or chronic disease that would preclude study participation; (6) having undergone body shape altering operation such as liposuction or plastic surgery. This data collection was approved by the Institutional Review Board (IRB).
Table 1: Demographic characteristics of study population.
Demographics N=60
Age(yr) $31.88\pm 6.04$
@@
   white 24
   African American 31
   others 5
Height(m) $1.62\pm 0.06$
Weight(kg) $71.91\pm 13.82$
Gestational age(d) $148.02\pm 8.84$
00footnotetext: Notes: For terms with continuous values, we provide mean±$\pm$±SD; for terms with discrete values, we provide number of participants.
@@
EFW(g) $378.40\pm 71.92$ (N=43)
MVP(cm) $4.90\pm 0.97$ (N=40)
Length of gestation(d) $268.69\pm 17.33$ (N=42)
Participants with cesarean section delivery 14 (N=42)
Participants with preterm labor 7 (N=42)
Participants with GDM 8 (N=51)
Participants with GH 8 (N=60)
00footnotetext: Notes: $N$ denotes the number of participants who have corresponding medical records. Since we have 6 participants transferred to other hospitals, 12 not delivered yet, 17 participants not having EFW value, 20 not having MVP value on their report, the number $N$ varies in different tasks. For terms including EFW, MVP and length of gestation, we provide mean±$\pm$±SD; for terms with binary results, we provide number of positive participants.
In addition to 3D scan data, we collected corresponding gestational age (GA) and basic demographic information including height and weight. To explore which adverse pregnancy outcomes and parameters are associated with body shape, we also collected the following information: EFW, MVP, indications of GDM and GH, length of gestation, and delivery type. The detail of these clinical parameters are listed in Table 2. Anthropometric measurements automatically generated by Fit3D scanner are used as the input for a SVM algorithm, which serves as the baseline method for our experiments. For each task, subjects who had incomplete information or those whose 3D scans were all unsuccessful were excluded from subsequent analysis.
@@ 4.1 Sequence-dependent feature extraction stream
Unlike a typical neural network unit, the RNN unit processes not only the current input, but also considers the output from the previous operation. As depicted in Figure 4, the RNN unit processing the element with index $i$ uses the current level circumference value $x_{i}$ and the previous output $h_{i-1}$ as inputs. The two inputs are given different trainable weights $W_{x}$ and $W_{h}$. After adding a trainable bias $b$, the weighted sum is processed through a nonlinear activation function $f(*)$. In our network, we specifically use the hyperbolic tangent function $tanh(*)$ to implement this nonlinear transformation. The output $h_{i}$ is then sent back to the RNN cell in the next step. The entire computation process can be formulated as follows:
@@ 4.3 Joint processing layer
Notice here both the dimensions of $vec_{PCA}$ and $vec_{basic}$ are 3. While processing concatenated features with fully connected layer, the dimension of each feature can also be considered as a “weight” [45]. Thus, we set the dimension of $h_{out}$ as 5, which is closer to the lower dimension of other features but still able to contain enough information, so that none of these three features can dominantly contribute to the output. In our experiments, we make minor adjustment on the final layer for different tasks. For regression tasks, we directly use this $output$ value as the final result and use the Mean Square Error (MSE) loss as criterion. For binary classification tasks, we generate a probability ranging from 0 to1 by using sigmoid function on the $output$ and use corresponding Binary Cross-Entropy (BCE) loss as criterion.
@@ 5 Experiments
We conducted binary classification tasks using 3D body shape features for the prediction of various adverse pregnancy outcomes, including the risk of preterm labor, GDM, GH, and the likelihood of undergoing a cesarean section during delivery. We also conducted regression tasks to estimate current fetal weight and MVP with the same inputs.
To assess the effectiveness of employing 3D body shape in these tasks and evaluate the performances of different algorithms, we implemented different evaluation metrics for classification tasks and regression tasks. Additionally, we designed an ablation study to validate the necessity and impact of critical components of our algorithm.
@@ 5.2 Classification Performance
We conducted binary classification experiments to predict the risks of preterm labor, GDM, GH, and the likelihood of undergoing a cesarean section. To evaluate the effectiveness of using 3D body scans compared to traditional anthropometric measurements, we developed a baseline method using anthropometric measurements automatically generated by the optical scanner (height, weight, BMI, waist circumference, hip circumference, waist-to-hip ratio, waist-to-height ratio) and GA as inputs. Support Vector Classification (SVC) was used as the baseline model to process these inputs. We also conducted comparison experiments with several popular machine learning algorithms using the same 3D body shape and demographic features as our algorithm used. These algorithms include Logistic Regression (LoR), BP Neural Network (BPNN), Random Forest (RF), and SVC. Since these algorithms cannot effectively process high dimensional features given small data size, we incorporate PCA to help improve their performance and mitigate potential overfitting problem. A 5-fold cross-validation was used in the experiments to ensure the algorithms’ generalizability and to provide more reliable comparisons.
@@
Baseline 42.86% 36.11% 92.86% 17.86% 0.52 0.531 (0.7458)
PCA+LoR 54.76% 35.29% 42.86% 60.71% 0.39 $<0.5$ (N/A)
PCA+BPNN 57.14% 30.00% 21.43% 75.00% 0.25 $<0.5$ (N/A)
PCA+RF 54.76% 36.84% 50.00% 57.14% 0.42 0.551 (0.5937)
PCA+SVC 42.86% 14.28% 14.28% 57.14% 0.14 $<0.5$ (N/A)
Ours 28.57% 25.00% 57.14% 14.28% 0.35 $<0.5$ (N/A)
@@
Baseline 80.95% 33.33% 14.28% 94.28% 0.20 0.592 (0.4486)
PCA+LoR 71.42% 22.22% 28.57% 80.00% 0.25 $<0.5$ (N/A)
PCA+BPNN 78.57% 37.5% 42.86% 85.71% 0.40 $<0.5$ (N/A)
PCA+RF 69.05% 25.00% 42.86% 62.86% 0.26 0.594 (0.4370)
PCA+SVC 83.33% 50.00% 28.57% 94.28% 0.36 0.539 (0.7471)
Ours 88.10% 75.00% 42.86% 97.14% 0.55 0.682 (0.1323)
@@
Baseline 68.00% 21.43% 37.50% 73.81% 0.27 0.613 (0.3140)
PCA+LoR 76.47% 30.00% 37.50% 83.72% 0.33 0.703 (0.0705)
PCA+BPNN 84.31% 50.00% 37.50% 93.02% 0.43 0.743 (0.0304)
PCA+RF 86.27% 66.67% 25.00% 97.67% 0.36 0.695 (0.0823)
PCA+SVC 84.31% 50.00% 37.50% 93.02% 0.43 0.637 (0.2222)
Ours 88.24% 75.00% 37.50% 97.67% 0.50 0.843 (0.0022)
@@
Table 6: Performance of classification methods for GH prediction
@@
Baseline 78.33% 14.29% 12.50% 88.46% 0.13 $<0.5$ (N/A)
PCA+LoR 83.33% 41.67% 62.50% 86.54% 0.50 0.827 (0.0031)
PCA+BPNN 85.00% 40.00% 25.00% 94.23% 0.31 0.669 (0.1263)
PCA+RF 83.33% 41.67% 62.50% 86.54% 0.50 0.867 (0.0009)
PCA+SVC 88.33% 55.56% 62.50% 92.31% 0.56 0.776 (0.0125)
Ours 88.33% 60.00% 37.50% 96.15% 0.46 0.738 (0.0313)
@@
The results are shown in Table 6, 6, 6, 6. We considered the situations with adverse outcomes as the positive class. For instance, in the delivery type classification task, participants who underwent a cesarean section are labeled as positive.
Figure 5: Visualization of raw 64 level circumferences data. The colors and line styles are assigned according to classification results of our algorithm in different tasks.
From these results, we can conclude that 3D body shapes captured in the second trimester can effectively be used to predict risk of preterm labor, GDM and GH, which can achieve accuracy rates exceeding 88%. However, using this data cannot reliably predict the delivery type. Even with the most effective algorithm for this task, the Random Forest, the accuracy only reaches 54.76%, with a precision of 36.84%, recall rate of 50%, and AUC-ROC of 0.551. This could be due to the complex factors influencing delivery type such as previous delivery experiences and umbilical cord status, which were not included in this study.
@@
By comparing the performances of the baseline method (SVC using anthropometric measurements as input) and the SVC using 3D body shape data, we can conclude that the same algorithm SVC can yields better prediction performance with 3D body shape as inputs than with traditional anthropometric measurements. In comparison to other algorithms, our algorithm shows superior performance in predicting risk of preterm labor, GDM and GH. Especially in predicting the risk of preterm labor and GDM, our algorithm surpasses others across all metrics. In the GH prediction task, our algorithm performs best in accuracy, precision, and specificity, while LoR, RF and SVC achieve a higher recall score of 62.5%. For predicting both GDM and GH tasks, our algorithm achieves high AUC-ROC scores with p-values lower than 0.05. These results suggest that 3D body shape data has higher potential to be used in predicting adverse pregnancy outcomes than anthropometric measurements. And these results also verified the effectiveness of our design in exploring useful information from 3D body scans.
To provide a more intuitive understanding of the relationship between abdominal body shape and adverse pregnancy outcomes, in Figure 5, the raw 64 abdominal level circumferences of subjects are presented as curve lines, and each curve is classified by colors and line styles according to the classification results yielded by our algorithm. In these charts, curves of negative subjects are drawn with blue lines and those of positive subjects are drawn with red lines. The curves of correctly classified subjects are drawn with solid lines and those of misclassified subjects are drawn with dashed lines. By observing the color of lines, we notice that the data is imbalanced, especially for preterm labor, GDM and GH. Despite the imbalance and small size of the dataset, our method yields better results in predicting preterm labor, GDM, and GH. Most negative samples are correctly classified, with false positive rates below 5%. Additionally, the details of the graph reveal intriguing information about the association between adverse outcomes and body shape patterns. We find that positive samples of cesarean section delivery do not have distinctive distribution. This may be one reason why we cannot accurately predict the cesarean section with 3D body shape and limited demographic information. Preterm labor and GDM exhibit similar positive sample distributions, which more likely happens to pregnant women with larger body circumferences. While positive GH samples tend to have relatively larger gravid uterus but smaller hip-region circumferences. From the chart we also notice some sudden changes in circumference values. This may be caused by scanner error or posture changes of participants.

### v2

@@ Abstract
Monitoring maternal and fetal health during pregnancy is crucial for preventing adverse outcomes. While tests such as ultrasound scans offer high accuracy, they can be costly and inconvenient. Telehealth solutions and more accessible body shape information provide pregnant women with a convenient way to monitor their health. This study explores the potential of 3D body scan data, captured during the 18-24 gestational weeks, to predict adverse pregnancy outcomes and estimate clinical parameters. We developed a novel algorithm with two parallel streams which are used for extract body shape features: one for supervised learning to extract sequential abdominal level circumference information, and the other for unsupervised learning to extract global shape descriptors, alongside a branch incorporating shape-related demographic data. Our results demonstrated that 3D body shapes can support the prediction of preterm labor and gestational diabetes mellitus (GDM), as well as the estimation of fetal weight. Compared to other machine learning models, our algorithm achieved the best performance, with prediction accuracies exceeding 89% and fetal weight estimation accuracy of 72.22% within a 10% error margin, outperforming the conventional anthropometric measurements-based method by 18.18%.
@@
Pregnancy outcomes, 3D body scan, Machine learning, Telehealth
@@ 1 Introduction
Although telehealth offers convenience, it has disadvantage of lacking in laboratory tests [10, 11]. To address this issue, it is crucial to explore biomarkers which can be easily obtained through widely-available devices for health monitoring. The 3D body shape is an ideal data modality that can be collected in non-invasive ways, especially in the field of obstetrics, where distinct body shape changes occur on pregnant women. Meanwhile, previous studies have shown a strong correlation between maternal anthropometric measurements and the health status of mothers and fetuses [12, 13, 14]. Ay et al. found that maternal body mass index (BMI) during pregnancy positively correlates with the fetal weight. Maternal height, pre-pregnancy BMI, and gestational weight gain are also found to correlate with potential risks of having a small or large for gestational age child [15]. Risk of preeclampsia (PEC) and cesarean delivery have been associated with maternal weight, height and body circumferences [16, 17]. These promising results indicate the potential for using body shape features as additional resources in telehealth for pregnancy. Moreover, comparing to traditional anthropometric features, detailed 3D body scan data can provide more body shape information, such as the shape, volume and position of gravid uterus.
@@
Currently, 3D body scanning technology in the field of obstetrics and gynecology is still in its early stages. Rather than being employed as a novel data modality for direct diagnosis, 3D scan data have primarilu been used as visualization tool or a medium to extract conventional anthropometric features. For example, Dathan-Stumpf et al. used 3D body scans to obtain pelvic measurements for vaginal breech delivery assessment [20], while Gradl applied structure light to monitor maternal abdomen changes during pregnancy and quantified abdominal deformation to evaluate fetal kicking activity [21]. However, these applications do not fully exploit the rich information contained in 3D body scan. In contrast, studies in the field of nutrition have shown strong potential of applying machine learning algorithms for 3D body shapes analysis. For instance, Wang et al. designed a bi-channel network to detect hepatic steatosis using 3D body shape reconstructed from CT iso-surfaces [22]. Feng et al. proposed a hierarchical neural network to estimate body compositions including lean mass, fat mass, and bone mineral content using whole-body optical scans[23]. These studies highlight the feasibility of applying machine learning–based 3D body shape analysis to other medical domains, including obstetrics.
In this study, we aim to leverage machine learning–based 3D body shape analysis for maternal and fetal health status assessment. Our main contributions are summarized as follows:
@@
    We used 3D body scans and basic body shape-related demographic information to predict the risk of preterm labor, gestational diabetes mellitus (GDM), PEC, and the likelihood of delivery by cesarean section. We also used the same body shape features to estimate fetal weight and maximum vertical pocket (MVP). To the best of our knowledge, this is the first study that uses 3D body scans of pregnant women to assess maternal and fetal health status.
@@
    We conducted experiments comparing our algorithm using 3D body scans as inputs with the baseline method using Logistic Regression (LoR, for classification tasks), Linear Regression (LR, for regression tasks), and anthropomorphic measurements. Additionally, we also compared our algorithm with other well-performing machine learning algorithms using the same 3D body shape inputs. The experimental results not only indicate the feasibility of using 3D body scans to assess maternal and fetal health status, but also demonstrate the precision and efficiency of our new approach.
@@ 2.1 Association between anthropometric measurement and pregnancy health status
As the foundation for applying 3D body shape analysis in obstetrics, the relationships between maternal anthropometric measurements and various adverse outcomes have been extensively studied, with strong correlations identified. Boucher et al. conducted a study which revealed associations between the probability of cesarean delivery and anthropometric measurements such as weight, BMI, waist circumference and skinfold thickness [17]. Sina et al. found that increasing BMI, weight, waist circumference, and waist-to-height ratio were associated with an increased risk of GDM, with BMI and waist-to-height ratio showing stronger correlations [24]. Ebrahimi-Mameghani et al. concluded that early pregnancy BMI and waist circumference were associated with the risk of gestational hypertension (GH), PEC, and preterm labor [25].
@@
Anthropometric techniques have been widely used to capture simplified body shape features for a long time, but these methods cannot fully represent the whole body shape [33, 34]. Moreover, although these methods usually do not require complex equipment, they still need to be carried out by professionals.
@@
Thus, some researchers investigated the applications of 3D body scanning in obstetrics. Gradl conducted a preliminary study monitoring maternal abdomen changes up to two months postpartum and evaluated fetal kicking activity during late pregnancy [21]. Glinkowski et al. used 3D surface topography method to reveal postural changes of pregnant women and studied the relationship between spinal curvatures and low back pain during pregnancy [35]. Dathan-Stumpf et al. captured anthropometric parameters from 3D scan models of pregnant women to predict successful vaginal breech deliveries. The results showed that the prediction accuracy of the method using 3D scanning measurements is at least as good as Magnetic Resonance Imaging (MRI) diagnosis [20]. These studies further indicated the usability of 3D body scans in obstetrics. However, these approaches remain limited to visualization purposes or the extraction of a small set of anthropometric measurements from 3D models, rather than directly leveraging the 3D models themselves as a new imaging modality.
@@
To fully leverage information from the complicated representations of 3D models, which are often comprised of thousands or millions of polygons or 3D points [36], researchers are investigating more efficient body representations and methodologies and evaluating them on different medical applications. Xie et al. proposed to use frontal whole-body silhouettes for estimating body composition. Test results on estimating fat mass index and fat-free mass index supported the usability of this silhouette-based method [37]. Ng et al. proposed a method to estimate regional body composition by using regional circumferences, areas, and volumes obtained from 3D body models with a linear model [18]. Lu et al. used level circumferences to represent 3D body shapes. With the extracted level circumferences, they employed a Bayesian network to predict pixel-level body composition and body fat percentage for the 2D projection of 3D body [38]. Wang et al. constructed a body shape descriptor called Shape Map from body shape contour of CT scan slices and used it to predict whole-body fat percentage and visceral fat percentage [39]. Roy et al. obtained the trajectory of the vertebral column from body shape and derived potential lateral deviations of the spine and rotation of the vertebrae. This method performed well in the estimations of the lateral deviation of the spine for mild and moderate scoliosis [40]. Su et al. explored the diagnosis of type II diabetes using features selected from 3D scan data through machine learning algorithms like backpropagation neural networks and decision trees [41]. Zheng et al. investigated the usability of raw 3D point cloud of body scans for deep learning-based body fat percentage prediction [42]. The same group also conducted a study comparing the effectiveness of different body shape representations, including abdominal level circumferences, area, and trunk 3D point cloud, in nonalcoholic fatty liver disease assessment [43]. Although these studies are not specifically conducted in obstetrics field, the various methods explored to represent 3D body shape and extract body shape features are worth learning.
@@ 3.1 Dataset
Our dataset comprises a total of 144 study participants, who were recruited at the George Washington University Medical Faculty Associates. The demographic characteristics of this study population is listed in Table 1. To mitigate the influence of confounding factors, we implemented the recruitment following these exclusion criteria: (1) being under the age of 18; (2) having carried multiple gestations; (3) having been diagnosed with an enlarged fibroid uterus; (4) having a BMI exceeding 60; (5) having any unstable medical or emotional condition or chronic disease that would preclude study participation; (6) having undergone body shape altering operation such as liposuction or plastic surgery. This data collection was approved by the Institutional Review Board (IRB). Informed consent forms were collected from all participants.
[…]
Demographics N=144
Age(yr) $32.98\pm 5.51$
@@
   white 76
   African American 53
   others 15
Height(m) $1.63\pm 0.07$
Weight(kg) $72.13\pm 13.92$
Gestational age(d) $146\pm 7.92$
@@
EFW(g) $382.48\pm 67.90$ (N=120)
MVP(cm) $4.75\pm 0.95$ (N=112)
Length of gestation(d) $271.40\pm 15.83$ (N=131)
Participants with cesarean section delivery 34 (N=129)
Participants with preterm labor 15 (N=131)
Participants with GDM 15 (N=138)
Participants with PEC 13 (N=130)
In addition to 3D scan data, we collected gestational age (GA) as a temporal marker, along with basic body shape-related demographic information including height and weight. To explore which adverse pregnancy outcomes and parameters can be predicted by body shape features, we also collected the following information: EFW (obtained from ultrasound examination report using Hadlock’s formula), MVP, indications of GDM and PEC, length of gestation, and delivery type [45]. The detail of these clinical parameters are listed in Table 2. Anthropometric measurements used as inputs for our baseline models were automatically generated by the Fit3D scanner. For each task, subjects who had incomplete information were excluded from subsequent analysis.
@@ 4.1 Sequence-dependent feature extraction stream
Unlike a typical neural network unit, the RNN unit processes not only the current input, but also considers the output from the previous operation. As depicted in Figure 4, the RNN unit processing the element with index $i$ uses the current level circumference value $x_{i}$ and the previous output $h_{i-1}$ as inputs. The two inputs are given different trainable weights $W_{x}$ and $W_{h}$. After adding a trainable bias $b$, the weighted sum is processed through a nonlinear activation function f()f(*). In our network, we specifically use the hyperbolic tangent function tanh()tanh(*) to implement this nonlinear transformation. The output $h_{i}$ is then sent back to the RNN cell in the next step. The entire computation process can be formulated as follows:
@@ 4.3 Joint processing layer
Notice here both the dimensions of $vec_{PCA}$ and $vec_{basic}$ are 3. While processing concatenated features with fully connected layer, the dimension of each feature can also be considered as a “weight” [47]. Thus, we set the dimension of $h_{out}$ as 5, which is closer to the lower dimension of other features but still able to contain enough information, so that none of these three features can dominantly contribute to the output. In our experiments, we make minor adjustment on the final layer for different tasks. For regression tasks, we directly use this outputoutput value as the final result and use the Mean Square Error (MSE) loss as criterion. For binary classification tasks, we generate a probability ranging from 0 to1 by using sigmoid function on the outputoutput and use corresponding Binary Cross-Entropy (BCE) loss as criterion.
@@ 5 Experiments
We conducted binary classification tasks using 3D body shape features for the prediction of various adverse pregnancy outcomes, including the risk of preterm labor, GDM, PEC, and the likelihood of undergoing a cesarean section during delivery. We also conducted regression tasks to estimate current fetal weight and MVP with the same inputs. The data were split for cross-validation and independent test set in a 7:3 ratio. A 5-fold cross-validation with early stopping was employed for model tuning to mitigate overfitting and enhance the generalizability of the algorithms.
To assess the effectiveness of employing 3D body shape in these tasks and evaluate the performances of different algorithms, we implemented different evaluation metrics for classification tasks and regression tasks. Heatmaps were used to interpret which body parts contributed more to the classification results, and the Bland-Altman plots were employed to compare the regression performances [48]. Additionally, we designed an ablation study to validate the necessity and impact of critical components of our algorithm.
@@ 5.2 Classification Performance
Binary classification experiments were conducted to predict the risks of preterm labor, GDM, PEC, and the likelihood of undergoing a cesarean section. To evaluate the effectiveness of using 3D body scans compared to traditional anthropometric measurements, we developed a baseline method using anthropometric measurements automatically generated by the optical scanner (height, weight, BMI, waist circumference, hip circumference, waist-to-hip ratio, waist-to-height ratio) and GA as inputs. Logistic Regression (LoR) was used as the baseline model to process these inputs. We also conducted comparison experiments with several popular machine learning algorithms using the same 3D body shape and shape-related demographic features as our algorithm used. These algorithms include LoR, Backpropagation Neural Network (BPNN), Random Forest (RF), and Support Vector Classifier (SVC). Since these algorithms cannot effectively process high dimensional features given small data size, we incorporated PCA to help improve their performance and mitigate potential overfitting problem.
@@
Baseline 74.36% 50.00% 30.00% 89.66% 0.38 0.590 (0.4012)
PCA+LoR 71.79% 42.86% 30.00% 86.21% 0.35 <0.5<0.5 (N/A)
PCA+BPNN 69.23% 41.67% 50.00% 75.86% 0.45 0.638 (0.1980)
PCA+RF 53.85% 27.78% 50.00% 55.17% 0.36 0.534 (0.7511)
PCA+SVC 64.10% 33.33% 40.00% 72.41% 0.36 0.566 (0.5382)
Ours 79.49% 75.00% 30.00% 96.55% 0.43 0.641 (0.1885)
@@
Baseline 79.49% 28.57% 40.00% 85.29% 0.33 0.771 (0.0530)
PCA+LoR 89.74% 66.67% 40.00% 97.06% 0.50 0.588 (0.5297)
PCA+BPNN 87.18% 50.00% 20.00% 97.06% 0.29 0.659 (0.2562)
PCA+RF 69.23% 23.08% 60.00% 70.59% 0.33 0.647 (0.2938)
PCA+SVC 46.15% 13.64% 60.00% 44.12% 0.22 0.582 (0.5581)
Ours 89.74% 66.67% 40.00% 97.06% 0.50 0.676 (0.2088)
@@
Baseline 85.37% 40.00% 40.00% 91.67% 0.40 0.667 (0.2311)
PCA+LoR 85.37% 40.00% 40.00% 91.67% 0.40 0.683 (0.1894)
PCA+BPNN 87.80% 50.00% 20.00% 97.22% 0.29 0.600 (0.4733)
PCA+RF 87.80% 50.00% 20.00% 97.22% 0.29 0.556 (0.6880)
PCA+SVC 85.37% 40.00% 40.00% 91.67% 0.40 0.694 (0.1642)
Ours 92.68% 75.00% 60.00% 97.22% 0.67 0.800 (0.0314)
@@
Table 6: Performance of classification methods for PEC prediction
@@
Baseline 66.67% 15.38% 50.00% 68.57% 0.24 0.529 (0.8509)
PCA+LoR 76.92% 14.29% 25.00% 82.86% 0.18 0.557 (0.7118)
PCA+BPNN 84.62% 25.00% 25.00% 91.43% 0.25 0.586 (0.5773)
PCA+RF 89.74% 50.00% 25.00% 97.14% 0.33 0.632 (0.3923)
PCA+SVC 79.49% 16.67% 25.00% 85.71% 0.20 0.500 (1.0000)
Ours 84.62% 25.00% 25.00% 91.43% 0.25 0.625 (0.4179)
@@
The results on the test sets are shown in Table 6, 6, 6, 6, with the best performances highlighted in bold. We considered the situations with adverse outcomes as the positive class. For instance, in the delivery type classification task, participants who underwent a cesarean section are labeled as positive.
These results suggest that 3D body shapes captured in the second trimester can be effectively used to predict the risk of preterm labor and GDM, achieving accuracy rates exceeding 89%. However, its performance is not very good in predicting the delivery type and PEC, where all of the methods get an F1 score lower than 0.5 and AUC-ROC lower than 0.650. This may be attributed to the complex factors influencing delivery type and the risk of PEC, such as obstetric history and gestational disease history, which were not considered in this study.
@@
By comparing the performances of the baseline method (LoR using anthropometric measurements as input) with LoR using 3D body shape data for preterm labor and GDM prediction, we find that the same LoR model achieves better performance when using 3D body shapes rather than traditional anthropometric measurements. Compared with other algorithms, our proposed method shows superior performance in predicting the risk of preterm labor and GDM. Especially for GDM prediction, our algorithm outperforms all others across all metrics, with a p-value below 0.05. Even in the more challenging delivery type prediction task, our algorithm achieves the highest accuracy, precision, specificity, and AUC-ROC, although BPNN and RF achieve a higher recall score of 50.00%. These results suggest that 3D body shape data have greater potential than anthropometric measurements for predicting adverse pregnancy outcomes and validate the effectiveness of our approach in extracting useful information from 3D body scans.
To provide a more intuitive understanding of which regions of abdominal body contribute more to predict adverse pregnancy outcomes, Figure 5 shows the average hidden state values for the negative and positive classes separately. From left to right, the 64 time points correspond to 64 level circumferences measured from the lowest point of the pubic bone to the bottom of the breasts. Referring to Figure 2, level circumferences 1–20 approximately cover the hip region, and level circumferences 30–40 roughly cover the waist region. For all tasks, the lower hip, top of the pelvis, and upper abdominal regions appear to play the most important roles, as indicated by greater changes in the hidden states. Except in the GDM prediction task, our model assigns extra attention to the lower hip region. Furthermore, we observe that for tasks on which our model performs well, there are clear differences between the heatmaps of negative and positive samples. In contrast, for tasks with poorer performance, the heatmaps are more similar, suggesting that the RNN branch may not extract sufficiently informative features from the body shape for downstream classifications.
Figure 5: Heatmaps of average RNN hidden states when processing the 64 level circumferences for different tasks. The upper rows labeled ”0” correspond to negative-class samples, and the lower rows labeled ”1” correspond to positive-class samples.